A random sample of 30 women drank an average of 15 cups of coffee per week, with the sample standard deviation equals to 3 cups. A upper limit of an approximate confidence interval with 95% of the population avg cup drunk is?
Answer by stanbon(75887)

A random sample of 30 women drank an average of 15 cups of coffee per week, with the sample standard deviation equals to 3 cups. A upper limit of an approximate confidence interval with 95% of the population avg cup drunk is?
-------------------------
sample mean = 15
---
standard error = E = t*s/sqrt(n)
E = invT(97.5 with df = 29)(3/sqrt(30)] = 2.0452*3/sqrt(30) = 1.1202
---
95% CI: 15-1.1202 < u < 15+1.1202
95% CI: 13.880 < u < 16.120
-----------------------------
Cheers,
Stan H.